OpenStudy (anonymous):

a car is moving at 4.5m/s acceleates at 2.5 m/s^2. After 7s it will be moving how fast?

OpenStudy (anonymous):

4.5+2.5*7=17.5+4.5=22 m/s

OpenStudy (anonymous):

final velocity=initial velocity+(acceleration x time)